A sports car moving at constant velocity travels 110 m in 5.0 s .

1.)If it then brakes and comes to a stop in 4.3 s , what is the magnitude of its acceleration (assumed constant) in m/s2?

2.)Express the answer in terms of "g's," where g=9.80m/s2.

Express your answer using two significant figures.
Answer #1

here,

initial constant speed , u = distance travelled /time taken

u = 110 / 5 m/s

u = 22 m/s

1)

let the accelration be a

time interval , t = 4.3 s

using first equation of motion

v= u + a * t

0 = 22 + a * 4.3

solving for a

a = - 5.12 m/s^2

the magnitude of acceleration is 5.12 m/s^2

2)

the magnitude of acceleration , |a| = 5.12 * g /9.80

|a| = 0.52 g